So I know this should be extremely easy, but for some reason I am brain dead. Simple question, need to know how to solve.
Find the arithmetic sum 2+5+8+...+101.
Thanks!
Jessica

Answer by stanbon(75887)

Find the arithmetic sum 2+5+8+...+101.
----
a(1) = 2
d = 3
--
Find the number of the term that equals 101:
2 + (n-1)3 = 101
3n-3 = 99
3n = 102
n = 34
----
Find the sum:
S(n) = (n/2)(1st + last)
S(34) = 17(2+101)
S(34) = 17(103) = 1751
==========================
Cheers,
Stan H.